# An Extensible Software Transport Layer for GPU Networking

### v1

@@ 7. Discussion
UCCL works around many challenges to support extensible transport with efficient multipathing on existing RDMA NICs, at the cost of QP swapping overhead (though not high for ML collectives), control coalescing, and more. UCCL performance and control granularity would get further boosted if there were a better HW-SW interface in RDMA NICs. We would like to highlight a few points based on our experience of developing UCCL: (1) UC abstraction is powerful with segmentation and reassembly offloading, but using many UC QPs for multipathing incurs some degree of QP swapping overhead; instead, it should evolve into a multipath UC abstraction that allows the software to specify different flow entropies for different verbs over a single QP, just like how UCCL specifies different UDP ports for AF_XDP. (2) NIC HW should expose more congestion signals to SW, such as ECN marks and packet trimming information; these signals can be embedded into CQEs just like hardware timestamps.
@@ 8. Other Related Work
Google GPUDirect-TCPX (Almasry and Willem de Bruijn, Eric Dumazet, Kaiyuan Zhang, 2023) integrates GPUDirect into the kernel TCP stack by leveraging the Header-Data Split feature in certain non-RDMA NICs. Instead, UCCL targets both RDMA and non-RDMA NICs, and further supports efficient multipathing. C4 (Dong et al., 2024) does coarse-grained flow-level traffic planning and path selection in LLM training, without programmability for low-level RDMA transport components such as CC and loss recovery. MSCCL (Azure/msccl, 2024) supports customizing collective communication algorithms and could work together with UCCL. Lastly, UCCL is inspired by a line of work that targets extensibility for CPU applications, e.g., Google 1RMA (Singhvi et al., 2020), eRPC (Kalia et al., 2019), and RoGUE (Le et al., 2018) for extensible RDMA transport, and SPIN (Bershad et al., 1995), Exokernel (Engler et al., 1995), and VINO (Seltzer et al., 1996) for extensible OS.
@@ 9. Conclusion
UCCL is an extensible and efficient software transport layer for GPU networking. It achieves network extensibility by separating the control path and data path for existing RDMA NICs and running the transport control path in software. Meanwhile, it achieves hardware-level performance by leveraging techniques like control coalescing and connection splitting. We hope UCCL opens the door to the productization of new research proposals on network transports for ML workloads. We will open source UCCL.
@@ C.1.2. Multi-collectives
In multi-collectives, GPUs with the same local rank in each server form a collective group, and multiple groups conduct collectives in parallel. For example, multi-allreduce is used in ML workloads with intra-server Tensor Parallelism (TP) + inter-server Data Parallelism (DP). We evaluate multi-collective performance by setting the environment variable NCCL_TESTS_SPLIT_MASK=0x7 in NCCL-tests. Figure 15, 15, 15 and 15 compare the multi-allreduce, multi-all-to-all, multi-allgather and multi-reduce-scatter performance of UCCL and SRD on the EFA testbed. UCCL outperforms SRD by up to 1.54, 1.22, 1.46, and 1.44 for multi-allreduce, multi-all-to-all, multi-allgather and multi-reduce-scatter, respectively.
@@ C.4.3. PCIe overhead
This experiment aims to study the PCIe overhead introduced by UCCL’s multipath design. We measure the PCIe overhead on our CX testbed, while the AWS virtualization layer prevents us from accessing low-level PCIe metrics. We rerun the all-to-all experiments in §6.1 and quantify MMIO events and CPU-NIC PCIe traffic using pcm-pcie (Intel, 2025). Figure 21 and 22 show the results. As expected, UCCL incurs higher MMIO activity and PCIe bandwidth consumption than the vanilla NCCL atop CX-7. UCCL incurs more MMIO events because it posts more verbs: UCCL transfers a small 32KB chunk per verb, while the vanilla NCCL atop CX-7 transfers a transport-buffer-sized message (e.g., default 128KB) per verb. The extra PCIe traffic of UCCL primarily comes from NIC swapping QPs (i.e., fetching uncached QP contexts from CPU memory) as UCCL uses 256 UC/RC QPs for multipath. UCCL UC consumes slightly more bandwidth than RC because it implements reliability and selective retransmission in software on the CPU.
We conclude that the number of MMIO events and extra PCIe traffic increase significantly with smaller chunk sizes, but the number of QPs does not nearly affect them. UCCL makes a reasonable trade-off between the control decision granularity and performance based on this observation and uses a 32KB chunk size as default (§3.3). We note that the extra PCIe bandwidth overhead is minor compared to the PCIe link capacity. For example, the capacity of PCIe 5.0 16 on CX is 512 Gbps in each direction, and the overhead of UCCL (32KB, 60k QPs) accounts for less than 2.5% (12/512=2.3%).

### v2

@@ 7. Discussion
HW-SW interface. UCCL works around many challenges to support extensible transport with efficient multipathing on existing RDMA NICs, at the cost of QP swapping overhead (though not high for ML collectives), control coalescing, and more. UCCL performance and control granularity would get further boosted if there were a better HW-SW interface in RDMA NICs. We would like to highlight a few points based on our experience of developing UCCL: (1) UC abstraction is powerful with segmentation and reassembly offloading, but using many UC QPs for multipathing incurs some degree of QP swapping overhead; instead, it should evolve into a multipath UC abstraction that allows the software to specify different flow entropies for different verbs over a single QP, just like how UCCL specifies different UDP ports for AF_XDP. (2) NIC HW should expose more congestion signals to SW, such as ECN marks and packet trimming information; these signals can be embedded into CQEs just like hardware timestamps.
GPU-driven communication like DeepEP (DeepSeek, 2025) leverages NVIDIA IBGDA (NVIDIA Corporation, 2025c) to issue RDMA verbs directly from GPUs to the RDMA NICs. Although seemingly conflicted with UCCL’s CPU-driven designs, GPU-driven communication could still be made compatible with UCCL. The key is to leverage the IBGDA CPU-assisted mode (NVIDIA Corporation, 2024b), where the GPU forwards RDMA requests to a CPU proxy that issues RDMA verbs; with this, we could implement UCCL’s extensible transport layer inside the CPU proxy. The trade-off is on performance: NVIDIA has reported that this CPU-assisted IBGDA would sacrifice 10% performance compared to the traditional IBGDA (NVIDIA Corporation, 2024b). We intend to leave the integration and performance enhancement of CPU-assisted IBGDA + UCCL as future work.
@@ 8. Other Related Work
Google GPUDirect-TCPX (Almasry and Willem de Bruijn, Eric Dumazet, Kaiyuan Zhang, 2023) integrates GPUDirect into the kernel TCP stack by leveraging the Header-Data Split feature in certain non-RDMA NICs. Instead, UCCL targets both RDMA and non-RDMA NICs, and further supports efficient multipathing. C4 (Dong et al., 2024) does coarse-grained flow-level traffic planning and path selection in LLM training, without programmability for low-level RDMA transport components such as CC and loss recovery. MSCCL (Azure/msccl, 2024) supports customizing collective communication algorithms and could work together with UCCL. Lastly, UCCL is inspired by a line of work that targets extensibility for CPU applications, e.g., Google 1RMA (Singhvi et al., 2020), eRPC (Kalia et al., 2019), RoGUE (Le et al., 2018), and IO-TCP (Kim et al., 2023) for extensible RDMA transport, and SPIN (Bershad et al., 1995), Exokernel (Engler et al., 1995), and VINO (Seltzer et al., 1996) for extensible OS.
@@ 9. Conclusion
UCCL is an extensible and efficient software transport layer for GPU networking. It achieves network extensibility by separating the control path and data path for existing RDMA NICs and running the transport control path in software. Meanwhile, it achieves hardware-level performance by leveraging techniques like control coalescing and connection splitting. We hope UCCL opens the door to the productization of new research proposals on network transports for ML workloads. UCCL is open-sourced at https://github.com/uccl-project/uccl.
@@ C.1.2. Multi-collectives
In multi-collectives, GPUs with the same local rank in each server form a collective group, and multiple groups conduct collectives in parallel. For example, multi-allreduce is used in ML workloads with intra-server Tensor Parallelism (TP) + inter-server Data Parallelism (DP). We evaluate multi-collective performance by setting the environment variable NCCL_TESTS_SPLIT_MASK=0x7 in NCCL-tests. Figure 18, 18, 18 and 18 compare the multi-allreduce, multi-all-to-all, multi-allgather and multi-reduce-scatter performance of UCCL and SRD on the EFA testbed. UCCL outperforms SRD by up to 1.54, 1.22, 1.46, and 1.44 for multi-allreduce, multi-all-to-all, multi-allgather and multi-reduce-scatter, respectively.
@@ C.4.3. PCIe overhead
This experiment aims to study the PCIe overhead introduced by UCCL’s multipath design. We measure the PCIe overhead on our CX_IB testbed, while the AWS virtualization layer prevents us from accessing low-level PCIe metrics. We rerun the all-to-all experiments in §6.1 and quantify MMIO events and CPU-NIC PCIe traffic using pcm-pcie (Intel, 2025). Figure 24 and 25 show the results. As expected, UCCL incurs higher MMIO activity and PCIe bandwidth consumption than the vanilla NCCL atop CX-7. UCCL incurs more MMIO events because it posts more verbs: UCCL transfers a small 32KB chunk per verb, while the vanilla NCCL atop CX-7 transfers a transport-buffer-sized message (e.g., default 128KB) per verb. The extra PCIe traffic of UCCL primarily comes from NIC swapping QPs (i.e., fetching uncached QP contexts from CPU memory) as UCCL uses 256 UC/RC QPs for multipath. UCCL UC consumes slightly more bandwidth than RC because it implements reliability and selective retransmission in software on the CPU.
We conclude that the number of MMIO events and extra PCIe traffic increase significantly with smaller chunk sizes, but the number of QPs does not nearly affect them. UCCL makes a reasonable trade-off between the control decision granularity and performance based on this observation and uses a 32KB chunk size as default (§3.3). We note that the extra PCIe bandwidth overhead is minor compared to the PCIe link capacity. For example, the capacity of PCIe 5.0 16 on CX_IB is 512 Gbps in each direction, and the overhead of UCCL (32KB, 60k QPs) accounts for less than 2.5% (12/512=2.3%).